### v4

# Interpretation of generalized Langevin equations

Abstract
Many real-world systems exhibit “noisy” evolution in time; interpreting their finitely-sampled behavior as arising from continuous-time processes (in the Itô or Stratonovich sense) has led to significant success in modeling and analysis in a wide variety of fields. Yet such interpretation hinges on a fundamental linear separation of randomness from determinism in the underlying dynamics. Here we propose some theoretical systems which resist easy and self-consistent interpretation into this well-defined class of equations, requiring an expansion of the interpretive framework. We argue that a wider class of stochastic differential equations, where evolution depends nonlinearly on a random or effectively-random quantity, may be consistently interpreted and in fact exhibit finite-time stochastic behavior in line with an equivalent Itô process, at which point many existing numerical and analytical techniques may be used.We put forward a method for this conversion, and demonstrate its use on both a toy system and on a system of direct physical relevance: the velocity of a meso-scale particle suspended in a turbulent fluid. This work enables the theoretical and numerical examination of a wide class of mathematical models which might otherwise be oversimplified due to a lack of appropriate tools.
I Generalizing Langevin Equations
Langevin equations are often used to represent theoretical differential behavior for systems exhibiting stochastic dynamics (see, e.g., [1, 2]). These equations have a standard form, which we will aim to generalize:
$\frac{\mathrm{d}x}{\mathrm{d}t}=f(x,t)+g(x,t)\eta_{t}\,,$
where $\eta_{t}$ represents the “Gaussian white noise” term, $\delta$-correlated in continuous time. If $g(x,t)$ exhibits $x$ dependence, such Langevin equations are ill-defined, necessitating a choice of either the Itô or Stratonovich interpretation—these will differ in the resulting “drift” behavior of the system, but both are internally consistent and able to be simulated and analyzed by various techniques [3, 2].
Here, we seek to generalize to systems of the form
$\frac{\mathrm{d}x}{\mathrm{d}t}=R(x,t),$ (1)
where $R$ is some random variable with some (possibly non-Gaussian) probability distribution over the domain. We argue that, with the proper conversion procedure based on the central limit theorem [4], these Langevin-type systems may be reduced to equivalent Itô behavior, allowing for consistent simulation and theoretical analysis.
As a motivating example, we start by highlighting the difference between two similar-looking Langevin-type equations:
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=-x^{3}+\eta_{t},$ (2)
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=-(x+\eta_{t})^{3}.$ (3)
Equation (2) is a classic Langevin equation with cubic attraction towards zero and diffusive noise—easily interpreted (in either the Itô or Stratonovich sense) as the stochastic differential equation (SDE) $\mathrm{d}x=-x^{3}\mathrm{d}t+\mathrm{d}W$ (where $\mathrm{d}W$ represents the usual derivative of a Wiener process), enabling all the analytical and numerical options that entails.
Equation (3), however, is notably different in that the nonlinear cubing operation happens to a fundamentally random quantity, linking the deterministic and random parts of the equation. Naïve numerical simulation simply converges to deterministic behavior as the time-step shrinks, since the fluctuations average out before $x$ changes considerably. If timestep-independent stochastic behavior is desired, we must develop a new consistent and coherent interpretation of this equation.
We note that the notation of Eq, (3) would be better written as
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=-X^{3},\textrm{ where }X\sim N(x,1).$ (4)
[…]
I.1 Potential Applications
“Baked-in” stochasticity of this type might arise in a variety of physical modeling scenarios. For example, nonlinear drag forces acting on a macroscopic object in a turbulent flow would cause velocity to evolve according to this type of Langevin equation, with “noise” coming from rapidly fluctuating relative fluid velocity—including, e.g., viscous drag on a cylinder in a turbulent wake [5]. We compute results for this velocity distribution, and its stark difference from a naïve approach, at the end of this section. Physical systems with nonlinear feedback based on rapidly fluctuating quantities or quantities subject to random measurement error would also be of this type. Inasmuch as measurement error acts as independent random variation of a quantity, the behavior of simulated or artificially-forced feedback-based dynamical systems would also benefit from this analysis. Our interest was motivated by an earlier model for individuals reacting to a stochastic political environment [6]. A variety of other physics-inspired nonlinear models of complex real-world phenomena may also share this form.
We note that the systems we are concerned with differ from other ways in which nonlinearity can arise in stochastic systems, for example in the deterministic part (e.g. [7]) or when $x$-dependence appears multiplied by the stochastic quantity (e.g., [8, 9]), or when functions are applied to a continuous random-walking quantity (as Itô’s lemma would handle [2]) rather than the uncertain/noisy quantity itself. Certain specific problems exhibiting nonlinear dependence on stochastic quantities have been examined [10], but a general theory of this class of stochastic equations has not been developed.
I.2 The Proposed Equivalence
We seek to bridge the gap from the theoretical, possibly non-Gaussian “intrinsic” noise (represented by the distribution $R$ in Eq. (1)) to some equivalent emergent system which is well-defined, self-consistent, and able to be simulated.
Our argument is based on the consideration that over any finite time-scale, a theoretical system such as Eq. (3) will have experienced a large enough number of nearly-independent increments that the Generalized Central Limit Theorem should apply [11]. That is, the net increment over any finite time must be drawn from the family of stable distributions, or—if the “intrinsic” noise represented by the differential update distribution itself has finite variance—a Gaussian distribution in particular [11]. This intuitively dovetails with the more practically-motivated necessary condition that, in the numerical simulation of any continuous-time system, its behavior must not depend sensitively on the simulated timestep; that is, one relatively large step must result in the same distribution (in an ensemble average sense) as the commensurate number of arbitrarily small steps.
We proceed henceforth with the assumption of finite underlying variance. This means that the net increment over any small but finite time must be drawn from a Gaussian distribution with mean equal to the mean of the underlying process. We may also choose this Gaussian’s distribution’s variance per unit time to likewise match the underlying process, maintaining consistency with the classic Langevin-Itô conversion and agreement in standard cases.
By this reasoning, we argue that every such stochastic process with finite variance is in fact equivalent to an Itô SDE over any finite time-scale: in particular, the SDE with deterministic part matching the “true” distribution’s mean behavior and random part matching its standard deviation. We note that this is not a one-to-one mapping, but rather many-to-one: any stochastic process with the same mean and standard deviation would behave identically, and thus be represented by the same Itô SDE.
That is, for a general stochastic system of the form
$\frac{\mathrm{d}x}{\mathrm{d}t}=R(x,t)\sim P(r|x,t),$
where $R$ is some finite-variance stochastic quantity dependent on $x$ and $\delta$-correlated in time, with distribution $P$, one should simulate the Itô SDE
$\qquad\mathrm{d}x=F(x,t)\mathrm{d}t+G(x,t)\mathrm{d}W,\quad\textrm{where}$
$F(x,t)=\textrm{mean}\left[R(x,t)\right]=\int\limits_{-\infty}^{\infty}rP(r|x,t)\mathrm{d}r,$
$G(x,t)=\textrm{std}\left[R(x,t)\right]=\sqrt{\int\limits_{-\infty}^{\infty}\left[r-F(x,t)\right]^{2}P(r|x,t)\mathrm{d}r},$
if these quantities exist. We will limit ourselves to stationary and autonomous processes (i.e., $F(x,t)=F(x)$ and $G(x,t)=G(x)$) for demonstration from this point forward, but the theory should extend to non-stationary processes.
Once we have this Itô equation, we may use standard numerical integration techniques for individual trajectories, or convert the system to a Fokker-Planck form and evolve the solution’s probability distribution $\rho(x)$ directly, with
$\frac{\partial\rho(x,t)}{\partial t}=-\frac{\partial}{\partial x}\left[F(x)\rho(x,t)\right]+\frac{1}{2}\frac{\partial^{2}}{\partial x^{2}}\left[G(x)^{2}\rho(x,t)\right]\;.$
As an example, we will now examine a slightly generalized version of Eq. (3) to determine the effect of noise with arbitrary constant amplitude $\sigma$:
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=-X^{3},\textrm{ where }X\sim N(x,\sigma).$ (5)
In section S1 of the Supplemental Material (SM), we examine a yet more general version of this attractor with arbitrary positive-integer exponent, but for illustration and concreteness henceforth focus on this cubic nonlinear-stochastic attractor. Using the shorthand notation $N(r|\mu,\sigma)=e^{\frac{-(r-\mu)^{2}}{2\sigma^{2}}}/(\sigma\sqrt{2\pi})$, we have:
$F(x|\sigma)$ $=\int\limits_{-\infty}^{\infty}-r^{3}N(r|x,\sigma)\mathrm{d}r$
$=-x^{3}-3\sigma^{2}x$
and
$G(x|\sigma)$ $=\sqrt{\int\limits_{-\infty}^{\infty}[-r^{3}-F(x)]^{2}N(r|x,\sigma)\mathrm{d}r}$
$=\sqrt{9\sigma^{2}x^{4}+36\sigma^{4}x^{2}+15\sigma^{6}}.$
So we argue that the system
$\frac{\mathrm{d}x}{\mathrm{d}t}=-X^{3},\textrm{ where }X\sim N(x,\sigma)$
is equivalent to the Itô SDE
$\mathrm{d}x$ $=(-x^{3}-3\sigma^{2}x)\ \mathrm{d}t$
$\qquad+\sqrt{15\sigma^{6}+36\sigma^{4}x^{2}+9\sigma^{2}x^{4}}\ \mathrm{d}W,$ (6)
which is amenable to various methods of simulation and analysis like any other Itô equation. We note that this Itô equation is significantly different from anything one might obtain from the similar-looking but simply additive true-Langevin form in Eq. (2).
To reiterate: a naïve interpretation of Eq. (5) (simply “expanding” $-(x+\sigma\eta_{t})^{3}$ and taking the pure-deterministic and pure-stochastic terms) might lead to the Itô SDE
$\mathrm{d}x=-x^{3}\mathrm{d}t+\sigma^{3}\mathrm{d}W\;,$ (7)
which has completely different physical behavior than our proposed interpretation in Eq. (6) [^1] In our “naïve” interpretation, we assume expansion of $(x+\sigma\eta)^{3}$ would be approximated with constant noise of amplitude $\sigma^{3}$, though the amplitude of constant noise could also be taken to be a fitted constant.. Basic properties like the variance of the equilibrium distribution differ, with divergence possible in Eq. (6) but not in Eq. (7). This has significant implications for all types of stochastic models used throughout physics.
I.3 Application to Drag in Turbulent Fluid
As an illustrative physical example, we consider the regime of quadratic drag with rapidly varying relative fluid velocity—of relevance to the behavior of particles in well-developed turbulence. In the one-dimensional case without stochasticity, relative velocity $v$ would evolve according to
$\textrm{d}v/\textrm{d}t=-cv|v|$
(here the constant $c$ sets the time scale, and we set it to 111 henceforth). When rapid random relative-velocity fluctuations are included, we have:
$\frac{\mathrm{d}v}{\mathrm{d}t}$ $=-V|V|\ ,\textrm{ \quad where }V\sim N(v,\sigma).$ (8)
This might naïvely be modeled by the Itô equation
$\mathrm{d}v=-v|v|\mathrm{d}t+\sigma^{2}\mathrm{d}W\;,$ (9)
which has an exact solution for its steady-state probability distribution
$p(v)=\frac{C}{\sigma^{2/3}}\exp{\left(\frac{-2|v^{3}|}{3\sigma^{2}}\right)}\;,$ (10)
where $C$ is a normalization constant, namely $3^{7/6}\Gamma(2/3)/(2^{5/3}\pi)$.
But we propose that this system is more faithfully modeled by using our proposed conversion, which yields
$\mathrm{d}v=F_{2}(v|\sigma)\mathrm{d}t+G_{2}(v|\sigma)\mathrm{d}W\;,$ (11)
where
$F_{2}(v|\sigma)$ $=-\left(\sigma^{2}+v^{2}\right)\operatorname{Erf}\left(\frac{v}{\sigma\sqrt{2}}\right)-\sqrt{\frac{2}{\pi}}x\sigma e^{\frac{-v^{2}}{2\sigma^{2}}}\;,$
$G_{2}(v|\sigma)$ $=\sqrt{v^{4}+6v^{2}\sigma^{2}+3\sigma^{4}+3[F_{2}(v|\sigma)]^{2}}$
(computation details in section S1 of the SM). The significant difference in behavior between these systems is illustrated in in Fig. 1.
Figure 1: Equilibrium velocity distributions. Comparison of equilibrium distributions for the drag system in Eq. (8) with $\sigma=0.2$, computed by Fokker-Planck integration of our proposed behavior (11) and compared to the exact solution (10) for a naïve interpretation of the system’s behavior.Top: Linear scale. Bottom: Log-scale view, with wider x domain to show clear differences in implied distributions.
II Discussion and Limitations
The first proposition of this paper—the argument for Ito-equivalency of nonlinear Langevin-type systems—is really a proposed definition rather than a theoretical result. Like Langevin equations themselves, the notation is simple and intuitive, but solid mathematical interpretation requires the use of the more rigorous notation, and we propose that interpretation in terms of Itô calculus—though this could easily be reinterpreted in the Stratonovich sense as convention or data favor.
We apply logic based on the central limit theorem for finite-variance random variables, but the Langevin noise terms are not regular random variables and their variance may not be well-defined or finite. If variance is treated as well-defined but not finite, other (non-Gaussian) stable distributions per time-step may arise, rather than normally distributed Itô time-steps.
We also note the perhaps-undesirable sensitivity to the assumption of Gaussian underlying noise in Eq. (1). In particular, the assumption that $\eta_{t}$ is normally distributed may be incorrect for some systems with biased or irregularly shaped noise, and if the noise shape is known it should be used.
III Conclusions
We have shown that a class of “nonlinear-stochastic” Langevin equations may be interpreted such that they have well-defined behavior after conversion to an equivalent Itô system. This type of equilibrium may be more general than initially apparent, since nearly any isolated attracting fixed point is locally well-approximated by equations of this form.
We have applied this theory to cubic and quadratic nonlinear attracting fixed points and arrived at appropriate equivalent Itô systems. These systems notably differ from any naïve system with $x$-independent noise, and Fokker-Planck numerical integration confirms that the implied equilibria of the quadratic case (a model of the velocity of a particle experiencing drag in a turbulent fluid) are quite different. In other work [21] we analyze the equilibria of the cubic system via moment relations and find that its higher moments diverge above a critical value of $\sigma$, and show that simulations bear out this analysis.
This conversion technique should lead to more faithful physical modeling, yielding qualitatively different behavior when compared to simplifications which transform a deterministic quantity and add noise afterward. Conversely, our reasoning also leads to the implication that apparent, observable Itô behavior might be emergent from any number of nonlinear Langevin processes at the theoretical level.
[…]
[width = 0.8color = black]88
Acknowledgements.
The authors thank Bill Kath for useful conversation, Gary Nave for help with relevant literature, and the National Science Foundation for support through the Graduate Research Fellowship Program.